If a rectangle has perimeter of 840ft and a length that is 188 feet more than the width, what is the length and width of the rectangle?
Answer by fractalier(6550)

Perimeter P = 2L + 2W = 840 and we also have L = W + 188. The substitute and we have
2(W + 188) + 2W = 840
2W + 376 + 2W = 840
4W + 376 = 840
4W = 464
W = 116 feet
L = W + 188 = 304 feet